OpenStudy (anonymous):

what is the conjugate of -2 + √(5)

OpenStudy (jhannybean):

A conjugate is the opposite binomial form of your function.

OpenStudy (jhannybean):

In this case, you would use the opposite operator in your function. main function: \(-2 + \sqrt{5}\) Conjugate: \(-2 \sqrt{5}\)

OpenStudy (nurali):

Conjugate of a complex number is simply a reversal of sign for the imaginary part. \[-2+\sqrt{5}\] Conjugate:\[-2-\sqrt{5}\]

OpenStudy (nurali):

When you multiply a complex number by it's conjugate, you get a real number.

OpenStudy (jhannybean):

oh did I make a typo...

OpenStudy (jhannybean):

Yes i did, lol.

OpenStudy (jhannybean):

I meant the conjugate to be \( -2 - \sqrt{5}\). Whoops.